# A Comprehensive Overview of Large Language Models

### 20 Feb 2024 v8

@@ VII Challenges and Future Directions
Overfitting: Although LLMs possess substantial learning capabilities, they are susceptible to overfitting noisy and peculiar patterns within their extensive training data. Consequently, this may cause them to generate illogical responses [468]. The debate about Memorization vs. Generalization in LLMs is about finding the right balance. Memorization allows the model to remember specific details from its training data, ensuring it can provide accurate answers to precise questions. However, generalization enables the model to make inferences and produce responses for inputs it hasn’t seen before, which is essential for handling various real-world tasks. Striking the right balance is the challenge: too much memorization can lead to overfitting, making the model inflexible and struggling with new inputs [469].
@@
Reasoning and Planning: Some reasoning and planning tasks, even as seemingly simple as common-sense planning, which humans find easy, remain well beyond the current capabilities of LLMs evaluated using an assessment framework. This isn’t entirely unexpected, considering that LLMs primarily generate text completions based on likelihood and offer no solid guarantees in terms of reasoning abilities [471].
Hallucinations: LLMs exhibit "hallucinations," where they generate responses that, while sounding plausible, are incorrect or don’t align with the provided information [472]. The hallucination can be categorized into three categories.
@@
Limited Knowledge: Information acquired during pretraining is limited and may become obsolete after some time. Re-training the model using updated data is costly. To generate factually accurate responses people use retrieval augmentation pipeline [187]. However, pre-trained models are not trained with retrieval augmentation generation (RAG) [6, 21], hence, adapting the training pipeline is necessary [182, 25].
@@
VIII Conclusion
This article has reviewed various LLMs, discussing their interesting aspects. It contributes to summarizing significant findings in the existing literature and provides a detailed analysis of the design aspects of LLMs, including architectures, datasets, and training pipelines. We identified crucial architectural components and training strategies employed by different LLMs. These aspects are presented as summaries and discussions throughout the article. Moreover, we have discussed the performance differences of LLMs in zero-shot and few-shot settings, explored the impact of fine-tuning, and compared supervised and generalized models and encoder vs decoder vs encoder-decoder architectures. A comprehensive review of LLMs in robotics, multi-modal LLMs, augmented LLMs, datasets, and evaluation is also provided. This article is anticipated to serve as a valuable resource for researchers, offering insights into the recent advancements in LLMs and providing fundamental concepts and details to develop better LLMs.

### 09 Apr 2024 v9

@@ 7 Challenges and Future Directions
Overfitting: Although LLMs possess substantial learning capabilities, they are susceptible to overfitting noisy and peculiar patterns within their extensive training data. Consequently, this may cause them to generate illogical responses [469]. The debate about Memorization vs. Generalization in LLMs is about finding the right balance. Memorization allows the model to remember specific details from its training data, ensuring it can provide accurate answers to precise questions. However, generalization enables the model to make inferences and produce responses for inputs it has not seen before, which is essential for handling various real-world tasks. Striking the right balance is the challenge: too much memorization can lead to overfitting, making the model inflexible and struggling with new inputs [470].
@@
Reasoning and Planning: Some reasoning and planning tasks, even as seemingly simple as common-sense planning, which humans find easy, remain well beyond the current capabilities of LLMs evaluated using an assessment framework. This is not entirely unexpected, considering that LLMs primarily generate text completions based on likelihood and offer no solid guarantees in terms of reasoning abilities [472].
Hallucinations: LLMs exhibit “hallucinations", where they generate responses that, while sounding plausible, are incorrect or do not align with the provided information [473]. The hallucination can be categorized into three categories.
  • 1.
@@
Limited Knowledge: Information acquired during pretraining is limited and may become obsolete after some time. Re-training the model using updated data is costly. To generate factually accurate responses people use a retrieval augmentation pipeline [188]. However, pre-trained models are not trained with retrieval augmentation generation (RAG) [6, 21], hence, adapting the training pipeline is necessary [183, 25].
@@
8 Conclusion
This article has reviewed the developments on LLMs comprehensively. It contributes to summarizing significant findings of LLMs in the existing literature and provides a detailed analysis of the design aspects, including architectures, datasets, and training pipelines. We identified crucial architectural components and training strategies employed by different LLMs. These aspects are presented as summaries and discussions throughout the article. Moreover, we have discussed the performance differences of LLMs in zero-shot and few-shot settings, explored the impact of fine-tuning, and compared supervised and generalized models and encoder vs. decoder vs. encoder-decoder architectures. A comprehensive review of multi-modal LLMs, retrieval augmented LLMs, LLMs-powered agents, efficient LLMs, datasets, evaluation, applications, and challenges is also provided. This article is anticipated to serve as a valuable resource for researchers, offering insights into the recent advancements in LLMs and providing fundamental concepts and details to develop better LLMs.